OpenStudy (anonymous):

last mounth 17 dogs and 2 cats were adopted estimate the probablyuty the next animal will be a cat write it in a percent

OpenStudy (anonymous):

If each month is independent of the last, the probability of adopting a cat is (number of cats)/(total number of animals) = 2/(17+2) = 2/19, which is about 10.5%.